# Multilevel Sampling in Algebraic Statistics

### v1

@@ 4.1 Maximum Mean Discrepancy
The Maximum Mean Discrepancy (MMD) [19] has been used over the past several decades as a non-parametric approach to determine if two distributions $P$ and $Q$ are different based on samples drawn from them. The MMD quantifies the difference between distributions using functions in a reproducing kernel Hilbert Space (RKHS). Specifically, it measures the discrepancy between two distributions by computing the difference between their mean embeddings in an RKHS. Given a positive definite kernel function $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$, the MMD is defined as
@@
The base kernel is often chosen as the distance-induced kernel given by $k(x,x^{\prime})=||x||_{2}+||x^{\prime}||_{2}-||x-x^{\prime}||_{2}$ for $x,x^{\prime}\in\mathbb{R}$. In this setting, the MMD reduces to the well known energy-distance [35].
@@ 4.2 Fiber Coverage Score
where $Z(0)=z_{0}$ is an initial feasible integer point in the fiber, $\mathcal{B}$ is a matrix whose columns form a lattice basis, $\sigma$ is a scale parameter, and $B(t)$ is a standard Brownian motion (see [31]). The sampler constructs real-valued combinations of lattice basis vectors using continuous Brownian paths, thereby generating points that lie within the polytope $\mathcal{P}_{A}(b)$. Points with negative entries (i.e., $Z(t)<0$ component-wise) are rejected and resampled. The resulting $Z(t)$ are real-valued vectors obeying the constraint $AZ(t)=b$, and serve as the centers of Voronoi cells. These are visualized as black dots in Figure 2. Once we obtain a set of centers $Z(t)$, we construct Voronoi cells around a selected subset of them and measure sample quality via the following volume-based metric.
@@ 5 Simulations
In the experiments that follow (excluding Experiment 2 in Section 5.1), we fix the polytope $\mathcal{P}_{A}(b)$ defined by the system $Ax=b$, with fiber $\mathcal{F}_{A}(b)$ drawn from a benchmark example in [3]; see Appendix A for full details. The following simulations evaluate the effectiveness of the methods introduced above, with a particular emphasis on the multilevel estimation algorithm presented in Section 3.1 (see also Appendix B for implementation details).
@@ 5.1 Multilevel Sampling
Figure 3: Multilevel estimate of the chi-square statistic distribution, obtained via Algorithm 1.
@@
This fiber was initially explored in Section 3, where we discovered issues arising from directly applying the procedure in Lemma 3.1. In order to address these issues, the multilevel sampler operates on both the larger and smaller scales, and obtains multiple Markov chains whose states (fiber points) come from both the local neighborhood around the initial point and the portions of the fiber further away from the initial fiber point. Results stand in stark contrast to the results from Figure 1 and thus we can already identify advantages of the multi-step approach. The chi-square distribution and the obtained sample cover a greater range of chi-square scores, giving us a good picture of the distribution of the tables in the fiber. In addition, there is no burn-in time in this approach, and the algorithm obtains the desired distribution in less iterations, as shown in Figure 5.
@@
For a concrete illustration, consider the infamous problem from [23]. The fiber $\mathcal{F}_{A}(b)$ in this case is a integer subset of polytope $\mathcal{P}_{A}$ defined by design matrix
@@
by sufficient statistic $b=\textbf{e}_{2k+1}$ and initial solution $x_{0}=(\textbf{0}_{k},\textbf{1}_{k},\textbf{0}_{k},\textbf{0}_{k},1,0)^{T}$. The parameter $k$ controls the size of the problem. Visually, the problem looks like Figure 4. The structure of the fiber is such that the fiber graph (pictured) has a single bottleneck edge, which poses a significant challenge for any randomized sampler. (The reader may be interested to know that [6] does sample this fiber successfully, after some careful parameter tuning.) However, the multilevel sampler encounters a different problem, namely that the $L_{2}$ norm of the initial solution $x_{0}=(\textbf{0}_{k},\textbf{1}_{k},\textbf{0}_{k},\textbf{0}_{k},1,0)^{T}$ is too small, and hence any levels larger than 1 produce unfeasible solutions with negative entries.
@@ 5.2 Sample Quality
To test whether multilevel strategies improve sample quality through broader exploration, we compare standard MCMC against our multilevel approach using the Maximum Mean Discrepancy (MMD) statistic defined in Equation (5). As a benchmark, we generate a large reference sample from the fiber of size $10^{5}$ using a standard MCMC algorithm with acceptance probability 1 and moves drawn from a Markov basis. This sample serves as a proxy for the true distribution over the fiber. Then, we compute the MMD between this benchmark and samples of size $n=500,1000,1500,\ldots,10000$ drawn independently from both the standard MCMC and multilevel methods. The multilevel sampler is implemented using the same experimental setup detailed in Appendix B. Each experiment is repeated over 30 trials, and the resulting mean MMD values with interquartile ranges are reported in Figure 5.
[…]
@@ 6 Discussion and Future Work
Several avenues remain open for future investigation. First, a key challenge lies in the optimal allocation of computational effort across levels. That is, how should one choose the number of samples $N_{\ell}$ at each level $\ell=1,\ldots,L$ in order to balance variance decay and cost? While theoretical guidelines exist in the continuous setting [18], analogous results in discrete spaces with Markov bases are undeveloped. Deriving adaptive or asymptotically optimal allocation rules for such settings is an important next step. Additionally, while our current estimator relies on equal-in-law subsampling between levels to ensure unbiasedness, more sophisticated coupling strategies could improve efficiency further. Finally, we again emphasize the importance of developing principled diagnostics for sample quality in discrete spaces. We introduced a Voronoi-based FCS quality metric and demonstrated its use in benchmarking different samplers. Further study of geometric or information-theoretic criteria for fiber exploration could provide new theoretical tools for evaluating and improving discrete sampling algorithms more broadly.
Acknowledgments
Fred Hickernell was instrumental in getting the conversation started with Henry Wynn. Research of first author supported by National Science Foundation (DMS Grant #2316011). Research of second and third authors partially supported by the Simons Foundation Gift # 854770 and DOE Grant #1010629.
@@ Appendix A Fiber Details
To this end, we consider a synthetic data set described in [12, p.371]. Note that their discussion is about $10\times 10$ tables, but we consider a $4\times 4$ sub-example on job-satisfaction, since it already proves our point. Our data is the same $4\times 4$ subset found in [3, Example 3.13]. The table classifies individuals according to $k=2$ categorical variables: $d_{1}=4$ levels of salary ranges and $d_{2}=4$ levels of job satisfaction/dissatisfaction. The table $U$ is thus of size $4\times 4$:
@@
For this example, there are $185,227,230$ tables in the fiber of the model of independence. Even if a lattice basis manages to explore the fiber (albeit at a slower rate, as the acceptance probability is usually lower than with Markov bases), the resulting $p$-values are not the same. Specifically, it appears that the lattice basis chain is missing a statistically significant part of the fiber, often not rejecting the model at the nominal level of $0.05$. This is illustrated in [3, Figures 1 and 2].
@@
$m=[1,-1,0,0,-1,1,0,0,0,0,0,0,0,0,0,0]$ satisfies $Am=0$, and is an element of a Markov and lattice bases for $A$. In particular, adding $m$ to the observed table $u$ does not change the margins of the table. Of course, this move alone cannot be applied to the starting data table, as it would create negative entries, thus stepping outside the set $\mathcal{F}_{A}(b)$.
The job satisfaction table example is a modified version of the following data on job satisfaction from [2, p.57]: $u=[1,2,1,0,3,3,6,1,10,10,14,9,6,7,12,11]$. The original job satisfaction data is also used in the documentation page for the R package algstat [26]. It is not too difficult to analyze, which is why we consider the synthetic data set for which lattice bases samplers already demonstrably do not perform well on average.

### v2

@@ 4.1 Maximum Mean Discrepancy
The Maximum Mean Discrepancy (MMD) [MMD2012] has been used over the past several decades as a non-parametric approach to determine if two distributions $P$ and $Q$ are different based on samples drawn from them. The MMD quantifies the difference between distributions using functions in a reproducing kernel Hilbert Space (RKHS). Specifically, it measures the discrepancy between two distributions by computing the difference between their mean embeddings in an RKHS. Given a positive definite kernel function $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$, the MMD is defined as
@@
[…]
@@ 4.2 Fiber Coverage Score
where $Z(0)=z_{0}$ is an initial feasible integer point in the fiber, $\mathcal{B}$ is a matrix whose columns form a lattice basis, $\sigma$ is a scale parameter, and $B(t)$ is a standard Brownian motion (see [Mörters_Peres_2010]). The sampler constructs real-valued combinations of lattice basis vectors using continuous Brownian paths, thereby generating points that lie within the polytope $\mathcal{P}_{A}(b)$. Points with negative entries (i.e., $Z(t)<0$ component-wise) are rejected and resampled. The resulting $Z(t)$ are real-valued vectors obeying the constraint $AZ(t)=b$, and serve as the centers of Voronoi cells. These are visualized as black dots in Figure 2. Once we obtain a set of centers $Z(t)$, we construct Voronoi cells around a selected subset of them and measure sample quality via the following volume-based metric.
@@ 5 Simulations
In the experiments that follow (excluding Experiment 2 in Section 5.1), we fix the polytope $\mathcal{P}_{A}(b)$ defined by the system $Ax=b$, with fiber $\mathcal{F}_{A}(b)$ drawn from a benchmark example in [MB25years]; see Appendix A for full details. The following simulations evaluate the effectiveness of the methods introduced above, with a particular emphasis on the multilevel estimation algorithm presented in Section 3.1 (see also Appendix B for implementation details).
@@ 5.1 Multilevel Sampling
Figure 3: Multilevel estimate of the chi-square statistic distribution, obtained via Algorithm 1. The underlying fiber sample is from the uniform distribution.
@@
This fiber was initially explored in Section 3, where we discovered issues arising from directly applying the procedure in Lemma 3.1. In order to address these issues, the multilevel sampler operates on both the larger and smaller scales, and obtains multiple Markov chains whose states (fiber points) come from both the local neighborhood around the initial point and the portions of the fiber further away from the initial fiber point. Results stand in stark contrast to the results from Figure 1 and thus we can already identify advantages of the multi-step approach. The chi-square distribution and the obtained sample cover a greater range of chi-square scores, giving us a good picture of the distribution of the tables in the fiber. In addition, there is no burn-in time in this approach, and the algorithm obtains the desired distribution in less iterations, as shown in Figure 5(a).
@@
For a concrete illustration, consider the infamous problem from [Hemmecke2014OnTC]. The fiber $\mathcal{F}_{A}(b)$ in this case is a integer subset of polytope $\mathcal{P}_{A}$ defined by design matrix
@@
by sufficient statistic $b=\textbf{e}_{2k+1}$ and initial solution $x_{0}=(\textbf{0}_{k},\textbf{1}_{k},\textbf{0}_{k},\textbf{0}_{k},1,0)^{T}$. The parameter $k$ controls the size of the problem. Visually, the problem looks like Figure 4. The structure of the fiber is such that the fiber graph (pictured) has a single bottleneck edge, which poses a significant challenge for any randomized sampler. (The reader may be interested to know that [RUMBA] does sample this fiber successfully, after some careful parameter tuning.) However, the multilevel sampler encounters a different problem, namely that the $L_{2}$ norm of the initial solution $x_{0}=(\textbf{0}_{k},\textbf{1}_{k},\textbf{0}_{k},\textbf{0}_{k},1,0)^{T}$ is too small, and hence any levels larger than 1 produce unfeasible solutions with negative entries.
@@ 5.2 Sample Quality
To test whether multilevel strategies improve sample quality through broader exploration, we compare standard MCMC against our multilevel approach using the Maximum Mean Discrepancy (MMD) statistic defined in Equation (5). As a benchmark, we generate a large reference sample from the fiber of size $10^{5}$ using a standard MCMC algorithm with acceptance probability 1 and moves drawn from a Markov basis. This sample serves as a proxy for a representative sample over the fiber. Then, we compute the MMD between this benchmark and samples of size $n=500,1000,1500,\ldots,10000$ drawn independently from both the standard MCMC and multilevel methods. The multilevel sampler is implemented using the same experimental setup detailed in Appendix B. Each experiment is repeated over 30 trials, and the resulting mean MMD values with interquartile ranges are reported in Figure 5(a).
Results show that the multilevel approach consistently achieves lower MMD values than standard MCMC for small sample sizes (e.g., $n=500\text{--}3000$), indicating improved approximation of the target distribution with fewer points. As $n$ increases, the performance gap narrows, and both methods converge to similar MMD values around $n=5000$, suggesting comparable approximation quality at larger sample budgets.
@@
To demonstrate the effectiveness of the multilevel algorithm on a different fiber and a model more complex than the independence model for two categorical variables, we repeat the procedure of Experiment 3, this time under the no-three-factor interaction model for three categorical variables [FienberBookCategorical]. This model postulates that the conditional relationship between any pair of variables given the third one is the same for all values of the third variable. The exact fiber specification is provided again in Appendix A.
The results in Figure 5(b) reinforce the conclusions from the previous experiment, and in fact highlight them more strongly. In this case, the fiber is substantially larger, making the advantages of MLMC more pronounced over MCMC. The multilevel exploration rapidly produces samples whose MMD statistic converges to the reference distribution, significantly outpacing the standard MCMC approach.
Experiment 5.
@@ 6 Discussion and Future Work
While we naturally assume availability of a Markov basis, we would be remiss not to include a discussion on the applicability of multilevel samplers in situations where a Markov basis is not available, as is the case for many problems in practice. We briefly mentioned in the introduction some literature on sampling without the entire basis, e.g. [HAT, DS98]. In fact, sampling fibers in the absence of Markov bases can be considered a research area on its own; we single out recent notable work [KahleYoshidaGarciaPuente] that combines Markov chains with sequential importance sampling, as well as incorporating reinforcement learning and Bayesian approaches with incomplete bases [RUMBA] and [RLFiberSampler]. In terms of machine learning, the fiber sampling problem using incomplete bases can be solved, in principle, using a progressive machine learning framework. Toward that end, [IvanPhD2025] introduces the idea of a progression on the coefficients of the linear combination of the lattice basis. Such a construction allows one to start with an easy-to-compute lattice basis and build more complex Markov moves using combinations, the idea being that one can control the size of the coefficients that are used in the linear combination of lattice basis vectors. Depending on the sparsity of the fiber and of the initial solution, a progressive sampler could start with smaller coefficients in a random combination of lattice basis elements, then incrementally increase the coefficients toward the theoretical upper bound. Although this is not a classical machine learning progression, this approach can be thought of as an analogue to the levels in the multilevel sampler presented in our paper. More work is required to fully investigate this direction, but initial simulations show promise that multilevel samplers can be used in this setup.
Several other avenues remain open for future investigation. First, a key challenge lies in the optimal allocation of computational effort across levels. That is, how should one choose the number of levels $L$, and the number of samples $N_{\ell}$ at each level $\ell=1,2,\ldots,L$ in order to balance variance decay and cost? While theoretical guidelines exist in the continuous setting [gil15_multi_density], analogous results in discrete spaces with Markov bases are undeveloped. Following the principles established in the multilevel Monte Carlo literature [Hei01a, giles2008, Gil15a], one selects $L$ so that the bias at the finest level is of the same order as the sampling error, ensuring that neither source of error dominates. In practice, this can be achieved by performing a short pilot run to estimate how the variance of the level differences decays with $\ell$, and then increasing $L$ until further refinements contribute negligibly to the variance relative to the desired tolerance. Thus, derivaing adaptive or asymptotically optimal allocation rules for algebraic statistics settings is an important next step where fibers may vary dramatically in size and structure: larger or more complex fibers may require deeper hierarchies to adequately control bias, whereas smaller fibers should be handled with fewer levels. Additionally, while our current estimator relies on equal-in-law subsampling between levels to ensure unbiasedness, more sophisticated coupling strategies could improve efficiency further. Finally, we again emphasize the importance of developing principled diagnostics for sample quality in discrete spaces. We introduced a Voronoi-based FCS quality metric and demonstrated its use in benchmarking different samplers. Further study of geometric or information-theoretic criteria for fiber exploration could provide new theoretical tools for evaluating and improving discrete sampling algorithms more broadly.
Acknowledgments.
Fred J. Hickernell was instrumental in getting the conversation started with Henry Wynn. Research of first author supported by National Science Foundation (DMS Grant #2316011). Research of second and third authors partially supported by the Simons Foundation Gift # 854770 and DOE Grant #1010629.
@@ Appendix A Fiber Details
We consider a synthetic data set described in [DS98, p.371]. Note that their discussion is about $10\times 10$ tables, but we consider a $4\times 4$ sub-example on job-satisfaction, since it already proves our point. Our data is the same $4\times 4$ subset found in [MB25years, Example 3.13]. The table classifies individuals according to $k=2$ categorical variables: $d_{1}=4$ levels of salary ranges and $d_{2}=4$ levels of job satisfaction/dissatisfaction. The table $U$ is thus of size $4\times 4$:
@@
For this example, there are $185,227,230$ tables in the fiber of the model of independence. Even if a lattice basis manages to explore the fiber (albeit at a slower rate, as the acceptance probability is usually lower than with Markov bases), the resulting $p$-values are not the same. Specifically, it appears that the lattice basis chain is missing a statistically significant part of the fiber, often not rejecting the model at the nominal level of $0.05$. This is illustrated in [MB25years, Figures 1 and 2].
@@
m=[1,1,0,0,1,1,0,0,0,0,0,0,0,0,0,0]m=[1,-1,0,0,-1,1,0,0,0,0,0,0,0,0,0,0] satisfies $Am=0$, and is an element of a Markov and lattice bases for $A$. In particular, adding $m$ to the observed table $u$ does not change the margins of the table. Of course, this move alone cannot be applied to the starting data table, as it would create negative entries, thus stepping outside the set $\mathcal{F}_{A}(b)$.
The job satisfaction table example is a modified version of the following data on job satisfaction from [Agresti2002, p.57]: $u=[1,2,1,0,3,3,6,1,10,10,14,9,6,7,12,11]$. The original job satisfaction data is also used in the documentation page for the R package algstat [algstat.R]. It is not too difficult to analyze, which is why we consider the synthetic data set for which lattice bases samplers already demonstrably do not perform well on average.
Experiment 4.
Consider the no-three-factor interaction model for three categorical random variables $X,Y,Z$ taking values $\{1,2\}\times\{1,2\}\times\{1,\dots,5\}$. For this model, also often called the model of homogeneous associations in the contingency tables literature, the toric matrix $A$ which computes the table marginals is the following:
A=[
1 0 0 0 1 0 0 0 1 0 0 0 1 0 0 0 1 0 0 0
0 1 0 0 0 1 0 0 0 1 0 0 0 1 0 0 0 1 0 0
0 0 1 0 0 0 1 0 0 0 1 0 0 0 1 0 0 0 1 0
0 0 0 1 0 0 0 1 0 0 0 1 0 0 0 1 0 0 0 1
1 1 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
0 0 1 1 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
0 0 0 0 1 1 0 0 0 0 0 0 0 0 0 0 0 0 0 0
0 0 0 0 0 0 1 1 0 0 0 0 0 0 0 0 0 0 0 0
0 0 0 0 0 0 0 0 1 1 0 0 0 0 0 0 0 0 0 0
0 0 0 0 0 0 0 0 0 0 1 1 0 0 0 0 0 0 0 0
0 0 0 0 0 0 0 0 0 0 0 0 1 1 0 0 0 0 0 0
0 0 0 0 0 0 0 0 0 0 0 0 0 0 1 1 0 0 0 0
0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 1 1 0 0
0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 1 1
1 0 1 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
0 0 0 0 1 0 1 0 0 0 0 0 0 0 0 0 0 0 0 0
0 0 0 0 0 0 0 0 1 0 1 0 0 0 0 0 0 0 0 0
0 0 0 0 0 0 0 0 0 0 0 0 1 0 1 0 0 0 0 0
0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 1 0 1 0
0 1 0 1 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
0 0 0 0 0 1 0 1 0 0 0 0 0 0 0 0 0 0 0 0
0 0 0 0 0 0 0 0 0 1 0 1 0 0 0 0 0 0 0 0
0 0 0 0 0 0 0 0 0 0 0 0 0 1 0 1 0 0 0 0
0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 1 0 1
[…]
For the experiment, we generated $570$ random triples of joint observations of $X,Y,Z$, which can be classified in the following $2\times 2\times 5$ contingency table:
[…]
=Z1=Z2=Z3=Z4=Z5
In the random data set, there were $39$ observations $\{X=1,Y=1,Z=1\}$, $27$ observations $\{X=1,Y=2,Z=1\}$, and so on. Flattened into a vector in $\mathbb{Z}^{20}$, the following is then the starting lattice point for the samplers:
u=[39,23,27,23,25,31,28,25,42,32,34,28,33,21,24,27,26,23,27,32].u=[39,23,27,23,25,31,28,25,42,32,34,28,33,21,24,27,26,23,27,32].
The Markov basis for this model consists of $10$ elements and can be found, for example, on the website [MBdatabase], or can be computed using a computer algebra software. This model, being the smallest non-decomposable model for contingency tables, is notorious for its Markov bases complexity [slimTables], which is unbounded as the number of states of the variables grows.